1. There is a put option for Euros with a strike price of $1.22 and a premium of $.06. There is another put option for Euros with a strike price of $1.16   and a premium of $0.03. You are slightly pessimistic about the Euro and decide to utilize a bear spread using these two put options.

a) Please draw the contingency graph and identify the maximum gain and loss as well as the breakeven point.

b) Assuming the resulting spot rate is $1.17 and each put option contract represents 50,000 Euros, what is the overall profit or loss?
Answer #1

Below are the Snippets from the excel of the Bear Spread using Put option.

Bear Strategy: Buying A Put option at the higher strike price and selling a Put option at the lower Strike price. Creating a limited gain from the bearish trend of the market while also limiting the loss at the bullish trend of the market.

a) As can be seen from the pay-off matrix at the second image that,

Max. Profit= $0.03

Max. Loss= $0.03

And Break-even point is at = $1.19 of market rate

b) If the Market/Spot rate is at $1.17 then,

Cash-flow from buy-side put option:

= Max(0,(Strike Price - Spot Price)) - Put Price

= $1.22 - $1.17 - $0.06

= $0.05 - $0.06

= -$0.01

Cash-flow from sell-side put option:

= -Max(0,(Spot Price - Strike Price)) + Put Price

= -Max(0,($1.17 - $1.16)) + $0.03

= $0.00 + $0.03

= $0.03

Adding above both cash flows, Profit/Loss at spot price of $1.17= -$0.01 + $0.03 = $0.02

BULLISH VERTICAL SPREAD WITH PUTS Buy Put and Sell Put at different strike prices BUY Put Jan Strike price $ Put price $ CashAT EXPIRY CASHFLOW Euro price PUT $1.22 PUT $1.16 Cash flow Profit/loss $ 1.10 $ 0.06 $ (0.03) $ 0.03 $ 0.03 $ 1.11 $ 0.05 $$0.08 $0.06 $0.04 - . - . - . - . - . - . - . - . - . - . - . - . - . - . - . - . $0.02 Sg $(0.02) $(0.04) $(0.06) $(0.08) $1